# The Procrustean Bed of Time Series: The Optimization Bias of Point-wise Loss

### v1

@@ 2.3 Non-linear Modeling of Non-deterministic EOB
To extend our analysis beyond linear systems, we now consider nonlinear, nondeterministic systems. A canonical approach for modeling such complexity is the Gaussian Mixture Model (GMM), representing the process distribution as a weighted sum of multiple Gaussian components:
Directly computing the non-deterministic EOB for a GMM is analytically intractable due to the logarithm-of-sums form, $\log\left(\sum_{k=1}^{K}\pi_{k}p(z_{k})\right)$, inherent in its log-likelihood. To overcome this obstacle, we reformulate the EOB using the language of information theory. By the Jensen’s Inequality, we have the lower bound of GMM non-deterministic EOB, detailed in Appendix E:
$\mathbb{E}[\mathcal{B}_{z}]\geq\sum_{k=1}^{K}\pi_{k}\mathbb{E}[\mathcal{B}_{k}]-H(\pi).$ (17)
@@ 3.1 Information-Theoretic Interpretation and Thinking
From an information-theoretic perspective, the non-deterministic EOB, $\mathbb{E}[\mathcal{B}_{z}]$, quantified in Theorem 2.6 and Eq. (17) is precisely the Kullback-Leibler (KL) divergence between the true joint distribution $P(z_{1:T})$ and the factorized distribution $Q(z_{1:T})=\prod_{t=1}^{T}p(z_{t})$ implicitly shielded by the point-wise independence assumption.
@@
This seemingly counter-intuitive result is the core of our analysis and aligns with the Paradigm Paradox (Theorem 2.2). $\mathbb{E}[\mathcal{B}_{z}]$ represents the irreducible lower bound of the total EOB, $\mathbb{E}[\mathcal{B}]$, stemming from the flawed point-wise independence assumption. This flaw is inherent to any point-wise loss function itself, irrespective of the model architecture being trained. Consequently, the more pronounced the internal structure of a sequence, the more critical temporal information is discarded. This leads to a powerful conclusion: data with the strongest structural signal is penalized most heavily by a structure-agnostic loss function.
This paradox offers a new theoretical lens to correct a classic empirical misdiagnosis: the difficulty of training simple models like MLPs on highly periodic data (Ziyin et al., 2020), stems not from limited model expressiveness, as often assumed, but from a fundamental optimization pathology. Our analysis confirms the problem is more fundamental, stemming not from the model, but from an optimization pathology. The inherent point-wise independence assumption actively works against learning the temporal structure of the data, fundamentally misaligned with the learning objective.
Critically, our findings reveal this optimization bias is fundamentally immune to the community’s current solutions: the bias cannot be eliminated by simply enhancing model capacity (e.g., using stronger networks (Wang et al., 2025c) or LLM (Jin et al., 2024b, c)) or employing more sophisticated optimization algorithms. As long as the training objective relies on the point-wise independence assumption, this inherent flaw will persist.
3.2 Moreover
More interesting discussions of the limitations and further work on EOB Theory are provided in Appendix F.
4 A Principled Debiasing Program
@@
Strategy 1: Statistical Alignment. Rather than modifying the temporal MSE loss function $\mathcal{L}_{\text{TMSE}}$ directly, a common heuristic approach to mitigate optimization bias is to introduce auxiliary terms that align local or global statistics $\mathcal{S}$. However, these adjusted loss functions do not address the root cause of EOB. Instead, they merely alleviate the independence issue of temporal MSE by imposing stronger external statistical constraints,
$\hat{\mathcal{L}}=\mathcal{\mathcal{L}}_{\text{TMSE}}+\mathcal{L}(\mathcal{S}).$ (19)
@@
$\hat{x}_{c}=\mathcal{T}_{\text{comp}}(x)\in\mathbb{R}^{T^{\prime}},\;x=\mathcal{T}_{\text{comp}}^{-1}(\hat{x}_{c})\in\mathbb{R}^{T},\;T^{\prime}<T.$ (20)
@@
5 A Concrete Solution via Fourier Transform
To instantiate the principles of the Debiasing Program, we select the Discrete Fourier Transform (DFT) as a canonical and highly effective implementation. The DFT is uniquely suited to this task as its key advantages: 1) approximate orthogonality decomposition and 2) exceptional computational efficiency with complexity $\mathcal{O}(L\log L)$. The comprehensive analysis of these advantages is provided in Appendix G.1. Furthermore, we demonstrate a rigorous gradient analysis of various DFT applications in Appendix H, yielding a genuinely effective optimization path.
Notably, global basis functions of DFT are ill-suited for decomposing non-stationary systems, while dynamic characteristics are widespread in practice to reduce SSNR. Discrete Wavelet Transform (DWT) emerges as a more practical alternative, leveraging its time-frequency localization capabilities and demonstrating superior performance in characterizing dynamic and localized patterns (See Appendix G.2).
5.1 Gradient Pathology Analysis of $\ell_{p}$ Norm
While the structural orthogonalization like DFT or DWT provides a theoretically sound solution, it introduces a new, practical challenge during training: the high dynamic range of the orthogonal elements, a scenario where standard $l_{p}$ norms exhibit critical optimization pathologies.
The $\ell_{2}$ norm suffers from gradient dominance: its gradient is proportional to the absolute error, causing high-magnitude components to overwhelm the learning process and leave weaker components under-optimized. Conversely, the $\ell_{1}$ norm’s constant-magnitude gradient leads to gradient fatigue: the uniform dynamic is too weak to drive large-magnitude components to convergence effectively, while may be too aggressive for small-magnitude components, causing overshooting. A detailed mathematical analysis of these gradient pathologies is provided in Appendix G.3.
@@ 5.2 Harmonized $\ell_{p}$ Norm Optimization
To resolve these duality of flaws, we propose the Harmonized $\ell_{p}$ Norm ($\mathcal{L}_{\text{Harm},\ell_{p}}$), a framework that dynamically re-weights the gradient for each component based on its historical magnitude, $\bar{f}_{k}$.
5.2.1 To Mitigate Dominance for $\ell_{2}$
Harmonized MSE (hMSE) up-weights the gradient of weaker components by incorporating a relative error term:
@@ 5.2.1 To Mitigate Dominance for $\ell_{2}$
5.2.2 To Cure Fatigue for $\ell_{1}$
Harmonized MAE (hMAE) amplifies the gradient for stronger components by scaling the loss to be proportional to each component’s magnitude:
@@ 5.2.2 To Cure Fatigue for $\ell_{1}$
5.2.3 Statistical Analysis of Harmonized Norm
We prove that the harmonized $\ell_{p}$ norm achieves both point-estimation and information-theoretic unbiasedness. Furthermore, it balances training stability and dynamic rectification (see Appendix G.4).
@@ 6.1 Experimental Settings
Benchmarks. The benchmark utilized in our work comprises two components: (1) A synthetic dataset combining a stochastic autoregressive series with a deterministic trigonometric series to validate the Paradigm Paradox (Theorem 2.2), and its generation mathematical mechanism is detailed in Appendix I.1. Specifically, the non-deterministic SSNR ($\textit{SSNR}_{z}=32$) and variance of innovation noise ($\sigma_{\epsilon}^{2}=0.25$) are fixed, and the total SSNR ($\textit{SSNR}_{x}$) is variable; (2) Eleven real-world benchmarks that cover various scenarios, each exhibiting unique characteristics and temporal resolutions, detailed in Appendix I.2. These are used to verify the proposed Principle Debiasing Program.
Backbones. Our baselines are categorized into three groups: (1) Transformer-based methods: iTransformer, PatchTST, Pyraformer, FEDformer, Autoformer; (2) MLP-based methods: TimeMixer, TSMixer, DLinear, FreTS; (3) CNN-based methods: TimesNet, MICN; (4) Others: FreDF. We provide detailed descriptions of backbones in Appendix I.3.
Implementation. The baseline models are reproduced using the scripts provided by (Liu et al., 2024) and (Wang et al., 2025b). Our $\mathcal{L}_{\text{Harm},\ell_{p}}$ is deployed on PatchTST model. Details are provided in Appendix I.4. Notably, we do not use the temporal MSE loss, relying solely on our $\mathcal{L}_{\text{Harm},\ell_{p}}$.
Figure 2: Error surfaces of Transformer with Gaussian distribution innovation. The blue and red arrows indicate the surface variation trend along horizon $h$ and the total SSNR $\textit{SSNR}_{x}$, respectively.
@@ 6.2 Empirical Verification of EOB Theory
Theoretical Optimality. To establish a rigorous baseline, we define the optimal MSE ($\text{MSE}^{\text{opt.}}$) as the variance of the purely stochastic component, $\sigma_{z}^{2}$, representing the irreducible error lower bound of an ideal predictor. Figure 2(c) visualizes the optimal relative MSE ($\text{MSE}_{\text{rel.}}^{\text{opt.}}$) surface which decays rapidly with increasing $\textit{SSNR}_{x}$,
$\text{MSE}^{\text{opt.}}$ $=\textit{SSNR}_{z}\cdot\sigma_{\epsilon}^{2}=\sigma_{z}^{2},$ (26)
$\text{MSE}_{\text{rel.}}^{\text{opt.}}$ $=\frac{\text{MSE}^{\text{opt.}}}{\sigma_{x}^{2}}=\frac{\textit{SSNR}_{z}}{\textit{SSNR}_{x}}.$ (27)
@@
(Inefficiency Ratio) For a well-trained model, the inefficiency ratio ($\eta$) is defined as the ratio of model’s actual MSE to the theoretically optimal MSE: $\eta=\frac{\text{MSE}_{\text{act.}}}{\text{MSE}^{\text{opt.}}}=\frac{\text{MSE}_{\text{act.}}}{\sigma_{z}^{2}}$ (28) where $\sigma_{z}^{2}$ is the variance of unpredictable stochastic part.
@@
0.374
0.305
0.374
0.317
0.384
0.335
0.397
0.334
0.414
0.333
0.407
0.313
0.392
0.264
0.353
0.271 0.364
0.316
0.397
0.366
0.433
Table 1: The Comprehensive Results on the Long-term Forecasting Task. Bold typeface and underlined text represent the champion and runner-up performance, respectively.
@@
96
0.381
0.389
0.382
0.399
0.376
0.394
0.393
0.408
0.379
0.389
0.382
0.400
0.412
0.421
0.412
0.428
0.390
0.403
0.396
0.411
192
0.430
0.417
0.437
0.430
0.428
0.421
0.447
0.440
0.427
0.419
0.430
0.433
0.450
0.439
0.457
0.453
0.444
0.438
0.446
0.442
336
0.485
0.446
0.485
0.452
0.479
0.441
0.489
0.463
0.464
0.438
0.474
0.460
0.508
0.472
0.506
0.479
0.480
0.456
0.488
0.466
720
0.495
0.466
0.518
0.488
0.476
0.475
0.514
0.499
0.460
0.460
0.529
0.508
0.510
0.489
0.523
0.498
0.494
0.495
0.512
0.510
Avg
0.448
0.430
0.455
0.443
0.440
0.433
0.461
0.453
0.433
0.426
0.454
0.450
0.470
0.455
0.475
0.464
0.452
0.448
0.461
0.457
ETTm1
96
0.313
0.342
0.322
0.360
0.327
0.353
0.349
0.379
0.326
0.353
0.327
0.367
0.327
0.359
0.334
0.375
0.340
0.360
0.355
0.379
192
0.365
0.368
0.365
0.385
0.378
0.376
0.388
0.397
0.373
0.377
0.371
0.391
0.375
0.385
0.389
0.403
0.381
0.380
0.389
0.396
336
0.394
0.390
0.395
0.407
0.420
0.408
0.424
0.420
0.407
0.398
0.399
0.408
0.409
0.409
0.417
0.423
0.412
0.402
0.419
0.418
720
0.458
0.429
0.458
0.444
0.490
0.448
0.494
0.459
0.465
0.432
0.460
0.446
0.491
0.454
0.490
0.461
0.472
0.439
0.477
0.454
Avg
0.383
0.382
0.385
0.399
0.404
0.396
0.414
0.414
0.393
0.390
0.390
0.403
0.401
0.402
0.407
0.416
0.401
0.395
0.410
0.412
Table 2: Plug-and-play Experiments of Harmonized $\ell_{1}$ Norm. Bold typeface denotes the champion performance.
Crucially, Figure 2(d) reveals that $\eta$ grows rapidly as $\textit{SSNR}_{x}$ increases, while rises and converges along $h$. This phenomenon strongly corroborates the insight of Paradigm Paradox: as the process becomes more deterministic (higher $\textit{SSNR}_{x}$) and theoretically “easier” to predict, the point-wise loss function incurs a progressively larger optimization bias. This forces the model to converge to a solution that is increasingly distant from the theoretical optimum. The elevated EOB signifies a fundamental conflict between the point-wise learning objective and the data’s intrinsic temporal structure. This phenomenon is consistent across diverse foundational deep learning architectures and data distributions (see Appendix J.1 for additional results), demonstrating the universality of our EOB theory irrespective of model expressiveness.

### v2

@@ 2.3 Non-linear Modeling of Non-deterministic EOB
To extend our analysis beyond linear systems, we now consider nonlinear, non-deterministic systems. A canonical approach for modeling such complexity is the Gaussian Mixture Model (GMM). We have the lower bound of GMM non-deterministic EOB via Jensen’s Inequlity:
$\mathbb{E}[\mathcal{B}_{z}]\geq\sum_{k=1}^{K}\pi_{k}\mathbb{E}[\mathcal{B}_{k}]-H(\pi)$ (16)
where $\pi_{k}$ is its corresponding mixture weight of component $k$ and $H(\pi)$ is the Shannon entropy of the mixture weights.
Proof. See Appendix E.
@@ 3.1 Information-Theoretic Interpretation and Thinking
From an information-theoretic perspective, the non-deterministic EOB, $\mathbb{E}[\mathcal{B}_{z}]$, quantified in Theorem 2.6 and Eq. (16), is precisely the Kullback-Leibler (KL) divergence between the true joint distribution $P(z_{1:T})$ and the factorized distribution $Q(z_{1:T})=\prod_{t=1}^{T}p(z_{t})$ implicitly imposed by the point-wise independence assumption.
@@
This seemingly counterintuitive result is the core of our analysis and aligns with the Paradigm Paradox (Theorem 2.2). $\mathbb{E}[\mathcal{B}_{z}]$ represents the irreducible lower bound of the total EOB, $\mathbb{E}[\mathcal{B}]$, stemming from the flawed point-wise independence assumption. This flaw is inherent in any point-wise loss function itself, irrespective of the model architecture being trained. Consequently, the more pronounced the internal structure of a sequence is, the more critical temporal information is discarded. This leads to a powerful conclusion: data with the strongest structural signal is penalized most heavily by a structure-agnostic loss function.
This paradox offers a new theoretical lens to correct a classic empirical misdiagnosis: the difficulty of training simple models on highly periodic data (Ziyin et al., 2020), stems not from limited model expressiveness, as often assumed, but from a fundamental optimization pathology. The inherent point-wise independence assumption actively works against learning the temporal structure of the data, fundamentally misaligned with the learning objective.
Critically, our findings reveal this optimization bias is fundamentally immune to the existing solutions: the bias cannot be eliminated by simply enhancing model capacity (e.g., using stronger networks (Wang et al., 2025c) or LLMs (Jin et al., 2024b, c)) or employing more sophisticated optimization algorithms. As long as the training objective relies on the i.i.d. assumption, this inherent flaw will persist.
Further Remarks.
More interesting discussions about the reinterpretation of time series analysis taxonomy and limitations on EOB Theory are provided in Appendix F.
4 Principled Debiasing Program
@@
Strategy 1: Statistical Alignment. Rather than modifying the temporal MSE loss function $\mathcal{L}_{\text{TMSE}}$ directly, a common heuristic approach to mitigate optimization bias is introducing auxiliary terms that align local or global statistics $\mathcal{S}$. However, these adjusted loss functions do not address the root cause of EOB. Instead, they merely alleviate the independence issue of temporal MSE by imposing stronger external statistical constraints,
$\hat{\mathcal{L}}=\mathcal{L}_{\text{TMSE}}+\mathcal{L}(\mathcal{S}).$ (18)
@@
$\hat{x}_{c}=\mathcal{T}_{\text{comp}}(x)\in\mathbb{R}^{T^{\prime}},$ $\quad x=\mathcal{T}_{\text{comp}}^{-1}(\hat{x}_{c})\in\mathbb{R}^{T}$ (19)
s.t. $T^{\prime}<T.$
@@
Notably, the use of a unitary transform alone is insufficient when the $\ell_{2}$ distance is considered (See Theorem H.1).
5 Concrete Solution of Debiasing Program
To instantiate the Principled Debiasing Program, we employ the Discrete Fourier Transform (DFT) as an implementation. The DFT is uniquely suited to this task as its key advantages: 1) approximate orthogonality decomposition and 2) exceptional computational complexity $\mathcal{O}(L\log L)$. The comprehensive analysis of these advantages is provided in Appendix G.1. Furthermore, we demonstrate a rigorous gradient analysis of various DFT applications in Appendix H.
Notably, global basis functions of DFT are ill-suited for decomposing non-stationary systems, while dynamic characteristics are widespread in practice. Discrete Wavelet Transform (DWT) emerges as a practical alternative in characterizing dynamic and localized patterns (See Appendix G.2).
@@ 5.1 Harmonized $\ell_{p}$ Norm Optimization
While the structural orthogonalization like DFT or DWT provides a theoretically sound solution, it introduces a new and practical challenge during training: the high dynamic range of the orthogonal elements, a scenario where standard $l_{p}$ norms exhibit critical optimization pathologies.
The $\ell_{2}$ norm suffers from gradient dominance. Conversely, the $\ell_{1}$ norm’s constant-magnitude gradient leads to gradient fatigue. A detailed mathematical analysis of these gradient pathologies is provided in Appendix G.3. To resolve these duality of flaws, we propose Harmonized $\ell_{p}$ Norm ($\mathcal{L}_{\text{Harm},\ell_{p}}$) which dynamically re-weights the gradient for each component via its historical magnitude, $\bar{f}_{k}$.
To Mitigate Dominance for $\ell_{2}$: Harmonized MSE (hMSE) up-weights the gradient of weaker components by incorporating a relative error term:
@@
To Cure Fatigue for $\ell_{1}$: Harmonized MAE (hMAE) amplifies the gradient for stronger components by scaling the loss to be proportional to each component’s magnitude:
@@
Statistical Analysis of Harmonized Norm: We prove that the harmonized $\ell_{p}$ norm achieves both point-estimation and information-theoretic unbiasedness. It balances training stability and dynamic rectification (See Appendix G.4).
@@ 6.1 Experimental Settings
Datasets. The datasets utilized in our work comprises two parts: (1) A synthetic dataset consisted of deterministic trigonometric series with stochastic autoregressive component with to validate the Paradigm Paradox, and its generation mathematical mechanism is detailed in Appendix I.1; (2) Eleven real-world datasets that cover various scenarios, each exhibiting unique characteristics and temporal resolutions to verify the proposed Principle Debiasing Program, detailed in Appendix I.2 .
Baselines. Our baselines are categorized into three groups: (1) Transformer-based methods: iTransformer, PatchTST, Pyraformer, FEDformer, Autoformer; (2) MLP-based methods: TimeMixer, TSMixer, DLinear, FreTS; (3) CNN-based methods: TimesNet, MICN; (4) Others: FreDF. We provide detailed descriptions in Appendix I.3.
Implementation. The baseline models are reproduced using the scripts provided by Liu et al. and Wang et al.. $\mathcal{L}_{\text{Harm},\ell_{p}}$ and FreDF are deployed on iTransformer. Details are provided in Appendix I.4. Notably, we do not use the temporal MSE loss, relying solely on our $\mathcal{L}_{\text{Harm},\ell_{p}}$ ($\gamma=0.5$, $\beta=0.3$).
@@ 6.2 Empirical Verification of EOB Theory
Figure 2: Error surfaces of Transformer with Gaussian distribution innovation. The blue and red arrows indicate the surface variation trend along horizon ($h$) and the total SSNR ($\textit{SSNR}_{x}$), respectively.
@@
Theoretical Optimality. To establish a rigorous baseline, we consider the theoretical $h$-step optimal forecast variance, denoted as $\text{MSE}^{\text{opt.}}$. While the exact value is determined by the accumulation of innovation variance (i.e., $\text{MSE}^{\text{opt.}}=\sigma_{\epsilon}^{2}\sum_{j=0}^{h-1}(h-j)\psi_{j}^{2}\leq\sigma_{z}^{2}$), for simplicity, we approximate the baseline using its asymptotic upper bound: $\text{MSE}^{\text{opt.}}\approx\sigma_{z}^{2}$. This serves as a conservative proxy for the irreducible error. Figure 2(c) visualizes the optimal relative MSE surface, which decays rapidly with increasing $\textit{SSNR}_{x}$,
$\text{MSE}_{\text{rel.}}^{\text{opt.}}$ $=\frac{\text{MSE}^{\text{opt.}}}{\sigma_{x}^{2}}\approx\frac{\sigma_{z}^{2}}{\sigma_{x}^{2}}=\frac{\textit{SSNR}_{z}}{\textit{SSNR}_{x}}.$ (25)
@@
(Inefficiency Ratio) For a well-trained model, the inefficiency ratio ($\eta$) is defined as the ratio of model’s actual MSE to the theoretically optimal MSE: $\eta=\frac{\text{MSE}_{\text{act.}}}{\text{MSE}^{\text{opt.}}}=\frac{\text{MSE}_{\text{rel.}}}{\text{MSE}^{\text{opt.}}_{\text{rel.}}}$ (26)
@@
0.390
0.375
0.448
Table 1: The Average of Comprehensive Results on the Long-term Forecasting Task. Bold typeface and underlined text represent the champion and runner-up performance, respectively.
@@
0.00183 0.03064
0.00173
0.02996
0.00277
0.03792
0.01333
0.08053
0.00317
0.03817
0.00444
0.04687
0.24795
0.36749
0.00225
0.03226
0.68197
0.37992
0.43848
0.41915
ETTh2
0.00267 0.03417
0.00266
0.03360
0.00436
0.04381
0.04715
0.13613
0.00536
0.04723
0.00477
0.05016
0.71687
0.56622
0.02046
0.10355
3.12829
1.31740
1.45494
0.84423
ECL
0.00045
0.01547
0.00071 0.01935
0.00197
0.03333
0.01022
0.08266
0.00302
0.03682
0.04169
0.14882
0.27091
0.38609
0.10550
0.23501
0.45887
0.41007
0.20771
0.29969
Exchange
0.00430
0.04105
0.01026
0.06610
0.00302 0.03636
0.01528
0.08504
0.00093
0.01268
0.01949
0.09404
0.02604
0.11808
0.03189
0.09171
0.60709
0.54045
0.80637
0.61896
Weather
0.00024
0.00897
0.00051 0.01331
0.00062
0.01573
0.00656
0.06103
0.00161
0.01332
0.00320
0.03868
0.24106
0.32620
0.00484
0.05016
0.40554
0.42633
0.13727
0.18382
BE
0.00337
0.03468
0.00414
0.03969
0.00545
0.04242
0.09990
0.17182
0.00517
0.00792
0.05758
0.02593
0.01368
0.04231
0.00125
0.01602
0.10126
0.15891
0.21130
0.26251
FR
0.00555
0.03635
0.00550
0.03803
0.00609
0.03728
0.10834
0.16371
0.00773
0.00760
0.04784
0.02656
0.01944
0.05884
0.00232
0.02983
0.09778
0.13491
0.17897
0.21835
Table 2: The Average of Comprehensive Results on the Missing Data Imputation Task. Bold typeface and underlined text represent the champion and runner-up performance, respectively.
Crucially, Figure 2(d) reveals that $\eta$ grows rapidly as $\textit{SSNR}_{x}$ increases, while rises and converges along $h$. This phenomenon strongly corroborates the insight of Paradigm Paradox: as the process becomes more deterministic (higher $\textit{SSNR}_{x}$) and theoretically “easier” to predict, the point-wise loss incurs a progressively larger optimization bias. This forces the model to converge to a solution that is increasingly distant from the theoretical optimum. The elevated EOB signifies a fundamental conflict between the point-wise learning objective and data’s intrinsic temporal structure. This phenomenon is consistent across diverse foundational deep learning architectures and data distributions (See Appendix J.1), demonstrating the universality of our EOB theory irrespective of model expressiveness.
6.3 Insight Experiments on Trigonometric Series
In response to the empirical misdiagnosis in Section 3.1, we evaluate the representation power of $\mathcal{L}_{\text{Harm},\ell_{p}}$ on synthesized pure trigonometric series. Figure 9 reveals a striking contrast: while the temporal MSE baseline captures the general trend, its predictions are plagued by high-frequency oscillations (Figure 9, left), reflecting a failure to cleanly capture the underlying periodicity. In the frequency domain (Figure 9, right), this failure is evidenced by a pervasive noise distribution across all frequency bands, suggesting that temporal MSE suffers from significant spectral corruption. In sharp contrast, the Transformer guided by $\mathcal{L}_{\text{Harm},\ell_{p}}$ achieves a near-perfect fit in both domains. These results provided strong empirical validation for the generality and effectiveness of our EOB theory. (See Appendix J.2)